40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14933     Entry time: Thu Oct 3 19:40:18 2019
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX/POY imbalance 

Summary:

There is an imbalance between the POX and POY detector outputs reported in the CDS system. Possibilities are (i) the POX PD has a uncoated glass window whereas POY does not or (ii) there is some problem in the elctronics.

Details:

  1. Nominally, we run the POX/POY locking with +18dB whitening gain on POY and +30 dB on POX. This is a factor of 4 difference.
  2. The DC levels reported in C1:LSC-POXDC_OUT and C1:LSC-POYDC_OUT differ by a factor of 10 (24 cts for POY vs 2.4 cts for POX with 0dB whitening gain). These channels come from the P2 connector on the back of the PD Interface board into the fast CDS system.
  3. The levels reported by the Acromag system (which come out of the P1 connector) are 60mV for POY  vs 15 mV for POX.
  4. I confirmed that this imbalance is not due to clipping on the POX photodiode - I tweaked the steering mirror and observed the plateau (I did not, however, look at the beam on the PD active area with an IR viewew which would be a more conclusive test).
  5. I measured the power incident on either PD (using Ophir power meter, filter OFF). They were both ~10uW, as expected since the beam extraction for POY and POX are identical - a single HR mirror and the vacuum viewport.

Update 820pm: 

  1. I checked that there is no glass window on the PD.
  2. It is hard to see the beam on a viewer - but with the PRM aligned, I think I convinced myself that the beam is pretty well centered on the PD. 

So increasingly, it looks like the electronics are the source of the problem.

ELOG V3.1.3-